OpenStudy (anonymous):

Using the Pythagorean theorem, find the hypotenuse (which is c) in cm: a= 160 cm b= 120 cm c= ?

OpenStudy (shane_b):

Use: \[a^2+b^2=c^2\] You know a and b...

OpenStudy (chaise):

Pythagoras theorem states that: \[a^2+b^2=c^2\] Does this help, or do you need further help?

OpenStudy (anonymous):

yes. i do

OpenStudy (shane_b):

go ahead @chaise

OpenStudy (anonymous):

so.. c= 200 cm or should c=200cm^2

OpenStudy (chaise):

a^2+b^2=c^2 Substituting in your values. 160^2+120^2=c^2 25600+14400=c^2 40,000=c^2 Taking the square root of both sides of the equal sign. c=200cm This is a length, and is only one dimensional, and thus has units of cm. If it was an area, it would have units cm^2, if it was a volume it could have units cm^3. You are right by saying 200cm.

OpenStudy (anonymous):

ok. Thanks